# Explainability of Text Processing and Retrieval Methods: A Survey

### v2

@@ Abstract.
Deep Learning and Machine Learning based models have become extremely popular in text processing and information retrieval. However, the non-linear structures present inside the networks make these models largely inscrutable. A significant body of research has focused on increasing the transparency of these models. This article provides a broad overview of research on the explainability and interpretability of natural language processing and information retrieval methods. More specifically, we survey approaches that have been applied to explain word embeddings, sequence modeling, attention modules, transformers, BERT, and document ranking. The concluding section suggests some possible directions for future research on this topic.
@@ 1. Introduction
Up to about 10 years ago, commonly used information processing tools generally extracted a feature vector — a list of features along with their associated numerical weights — from the form in which the information was initially provided: structured or unstructured text, images, audio, video, etc. These features typically corresponded to human-understandable properties or attributes of the original information. We confine our discussion in this article to textual information only. For text, features would include words and phrases occurring in the text, their counts, morphology, parts of speech, and various grammatical relationships. These features were chosen on the basis of linguistic considerations; they were, in a sense, interpretable by design. A number of studies (some of which are discussed in more detail in Section 6) did focus on understanding feature weights, their behaviour, and the mathematical / statistical models underlying their computation and use, but they were not explicitly designated as explainability / intepretability studies.
Over the past few years, advances in Deep Machine Learning (ML) techniques have revolutionised the state of the art in textual information processing (as well as a number of other areas). These improvements have come at a cost: text is no longer simply represented in terms of easily interpretable features. Instead, early in the information processing pipeline, text is converted to so-called “dense” vectors in $\mathbb{R}^{n}$; the dimensions of this space are no longer easily understood. This representation is then processed by ML systems that have large and complex architectures, involving many parameters, often numbering in the millions, or even billions. As a fallout, there has been a marked increase in the attention devoted to providing high-level, intuitive explanations of these models and methods.
Thus, this seems to be an appropriate time to present an organised summary of the major research efforts in the area of explainability in textual information processing and retrieval. Indeed, several such general surveys have already been published (Søgaard, 2021; Madsen et al., 2022; Zini and Awad, 2022; Anand et al., 2022), in addition to more specialised surveys, such as (Belinkov and Glass, 2019; Dufter et al., 2022). Our objective is to complement these existing surveys by describing more recent work in areas already covered there, and by focussing on areas not addressed by them. Before discussing the relation between this survey and the existing ones in Section 2, we first briefly describe how we collected the set of papers reviewed in this survey.
1.1. Scope of this survey
@@
The principal terms used in our search were variants of the keywords explanation (including explain*) and interpretability (viz., interpret*). These terms have been variously defined within the community. For example, according to (Ribeiro et al., 2016), explaining a model may be defined as presenting textual or visual cues that provide qualitative understanding of the relationship between the components of an input instance (e.g. words in text), and the model’s output for that instance. Doshi-Velez and Kim (2017) define interpretability more generally as simply “the ability to explain or to present in understandable terms to a human”. We adopt the broad sense conveyed by these definitions, and use the terms explainability and interpretability interchangeably in this survey.
@@
Our preliminary search revealed a substantial body of work on explainability in Recommender Systems and Computer Vision. However, we excluded papers from these areas unless they directly contributed to follow-up work in IR or NLP. While IR systems and Recommender Systems share similarities in their goal of ranking items to meet users’ information needs, often with or without explicit queries, this survey focuses specifically on the retrieval of text documents. As text-focused IR methods are deeply intertwined with NLP techniques, we concentrate on the explainability of IR and NLP to maintain this focus. We include papers published from 2004 to 2024 in this set. Some influential works were also added that were published long before 2004. From the longlisted set of papers, a subset of 45 papers was selected for discussion in this article.
@@
Organisation of the survey:  The remainder of this survey is organised as follows. Section 2 lists existing surveys that cover explainability in textual IR and NLP, and describes the relation between them and the current article. Section 3 introduces the basic terminology used in the literature on explainability, and presents a schema for classifying explainability research. The studies summarised in the rest of the article are broadly organised using this schema. Section 4 reviews some well-known approaches to explainability that are most commonly used as baselines in recent work. Section 6 is devoted to a detailed discussion of explainable document / passage ranking, and is itself divided into subsections by topic. Within each of these sections and subsections, different studies have been discussed in chronological order. Sections 7 through 10 are intended as supplements to existing surveys (specifically those by Madsen et al. (2022) and Zini and Awad (2022)); they cover recent work on explainability for individual elements of deep text processing pipelines, viz., word embeddings, sequence models, attention mechanisms, and transformer architectures. We conclude in Section 11 with a few research questions that we believe merit further investigation. 11todo: 1Update after everything is done.
@@ 2. Related Surveys
In a monograph (Lin et al., 2021), Lin et al. provide a comprehensive survey of neural IR models, covering topics from multi-stage ranking architectures to dense retrieval techniques. The authors observe that IR models may be stratified into three fairly well-separated levels on the basis of effectiveness: transformer-based ranking strategies represent a definite advance over pre-transformer Neural Ranking Models (NRMs), which in turn improve significantly over “traditional” [^1] In this survey, we use this adjective to refer to all non-neural IR models, including relatively recently proposed ones. statistical ranking models. An intuitive (but precise) grasp of what leads to these quantum jumps in effectiveness seems eminently desirable. This, along with a number of other factors, has led to much interest in the explainability of NRMs. There has always been a substantial overlap between the IR and NLP communities. With the recent success of neural methods, the boundary between the disciplines has been blurred even further. Not surprisingly, therefore, research in explainable IR (ExIR) is significantly influenced by similar research in NLP, which in turn draws heavily from explainability studies in ML in general.
@@
Our search for a comprehensive summary of research on this broad subject led us to two recent articles that provide an overview of attempts to explain neural models for NLP (Madsen et al., 2022; Zini and Awad, 2022). In (Madsen et al., 2022), the main emphasis is on post-hoc [^2] In Section 3, we provide a review of this and other basic terms commonly used in the context of interpretability / explainability. methods for interpreting neural NLP methods, while Zini and Awad (2022) focus on the explainability of word embedding vectors, recurrent neural networks (RNNs) and attention mechanisms.
@@
This being an area of active research, a non-trivial body of work exists that has not been covered by the above surveys, either because the work was published later, or because its topic was outside the specific scope of the surveys. In particular, until very recently, there were no similar overviews addressing ExIR in detail. The current study started as an attempt to address that lacuna. Recently, Anand et al. (2022) have published an article focused on this specific topic. There is, inevitably, a good deal of overlap between Anand et al.’s survey and Section 6 of this article, but we also cover a fair amount of work that is not discussed by them. Further, since our survey was prepared independently, we believe that it provides a somewhat different perspective on ExIR. Finally, given that ExIR often draws inspiration from explainability for NLP, we review (in Sections 7 through 10) recent work on Explainable NLP that was not covered by the earlier surveys (Madsen et al., 2022; Zini and Awad, 2022). In the rest of this section, we briefly discuss other related surveys that are either older, or have a scope that does not substantially overlap with ours.
@@
To the best of our knowledge, Doshi-Velez and Kim (2017) were the first to attempt to formalise various notions related to interpretability in ML. Lipton (2018) provides an exposition of the broader aspects of interpretability in ML and Artificial Intelligence (AI), and why it is challenging, especially in the modern scenario. Søgaard (2021) examines extant taxonomies of explainability methods, and carefully constructs a framework for organising them. The next section draws from these articles to build a glossary of common terms used in this area.
@@
Belinkov and Glass (2019) cover explainable NLP, and summarise studies that analyse neural network nodes, and activations across different layers of deep neural networks, in order to understand what linguistic features are encoded by neural NLP models. As mentioned above, many of the discussed methods were based on general interpretability approaches proposed by the ML community, often within the context of Computer Vision / Image Processing problems. Signalling the growing interest in this area, Danilevsky et al. (2020) also provide an early overview of explainability for NLP, describing different explanation categories and approaches, as well as several visualization techniques. Lertvittayakumjorn and Toni (2021) focus on explanation as a tool for debugging NLP models, with humans in the loop.
@@
As Transformer and BERT-based techniques were widely adopted in the IR/NLP community, several researchers focused on various aspects of BERT, from its knowledge-representation power to fine-tuning, pre-training, and over-parametrisation issues. Rogers et al. (2020) were the first to survey research in this category, and present a summary of “what we know about how BERT works.” The surveys discussed at the beginning of this section provide a more up-to-date summary of research in this broad area.
@@
There are other broader efforts in Explainable AI (XAI), such as XAIR (Clement et al., 2023), which systematically reviews XAI methods aligned with the software development process. However, this meta-review targets general AI explainability rather than the specific challenges of textual and retrieval systems addressed in our survey.
@@ 3. Categorising Explanation Methods
Some ML models, e.g., rule-based systems, decision trees and sparse linear models, are regarded as inherently explainable, because it is relatively simple for a human to trace and describe their inner workings. Traditional IR models like BM25 are also generally regarded as inherently explainable. Such models comprise the class $\mathcal{S}$ (see Table 1). Explanations of such models are, naturally, model-aware. The composition of $\mathcal{S}$ is open to debate, however. Moshkovitz et al. (2020) argue that even apparently well-understood methods with a single parameter, viz., $k$-means and $k$-medians clustering techniques, may be regarded as “opaque”, because these methods define cluster boundaries on the basis of linear combinations of features, rather than a single feature value. More pertinent to our context is the debate on whether the attention mechanism can be construed as providing explanations. Section 9.1 discusses this debate in more detail.
In contrast, post-hoc explanations primarily seek to elucidate the output, rather than the inner workings of models. Visualisations of word embeddings (e.g., using t-SNE (van der Maaten and Hinton, 2008)) are a form of post-hoc “explanation”: they do not explain why certain words are mapped to certain vectors, but once the vectors are generated (i.e., after the fact), they provide insights about the correspondence between proximity in the semantic and vector spaces. Explanations by example (e.g., “Model $M$ predicts class $Y$ for input instance $X$ because $X$ is ‘similar’ to training instances ${X}_{1},{X}_{2},\ldots,{X}_{k}$ which were labelled YY”) also fall within this category. Post-hoc explanations may be either model-aware or model-agnostic.
@@
Global methods explain the overall behavior of a model. Global explanations should apply to all (or most) input-output pairs for a given model. An example of a global explanation in the IR context is: “ranking based on query likelihood tends to be biased in favour of short documents when Jelinek-Mercer smoothing is used instead of Dirichlet smoothing” (Smucker and Allan, 2006). On the other hand, a local explanation only seeks to explain a model’s decision for a given input instance.
3.3.1. Pairwise vs. pointwise vs. listwise local explanations
@@ 3.4.2. Feature attribution and rationales
Instead of explaining $\mathcal{M}_{C}$’s behaviour in terms of $\mathcal{M}_{S}$, researchers have also attempted to understand $\mathcal{M}_{C}$ by identifying which features of an input instance $I$ (i.e., which components of $I_{S}$ or $I_{C}$) most strongly affect $\mathcal{M}_{C}$’s output for $I$. This approach is called feature attribution. Several techniques for quantifying and visualising the contribution of a particular feature towards the prediction made by a model have been proposed in the literature. These local explanations may be provided graphically in the form of a heatmap or saliency map (Simonyan et al., 2014), with different colours or shades representing different degrees of importance of features, or simply as a list of the most important features along with their importance scores.
A related approach involves the construction of a rationale, a concise representation (or the extracted essence) of $I$ that is (i) both short and simple enough to be readily understandable by a human, but (ii) for which $\mathcal{M}_{C}$’s prediction is the same as that for $I$. For example, a short excerpt from an input text that is sufficient to correctly predict the sentiment of a sentence comprises a rationale for a sentiment classification model (DeYoung et al., 2020).

### v3

@@ Abstract.
Deep Learning and Machine Learning based models have become extremely popular in text processing and information retrieval, owing to their remarkable effectiveness. However, the complex non-linear structures underlying these models make them largely inscrutable. A significant body of research has focused on increasing the transparency of these models. This article provides a broad overview of research on the explainability and interpretability of information retrieval methods, focusing primarily on both neural models for document ranking, and retrieval-augmented generation systems. A significant part of this research is inspired by, and strongly related to, similar work done in the area of natural language processing (NLP). For completeness, therefore, we also briefly review (in the Appendices) some recent studies from the NLP community on explaining word embeddings, sequence modeling, attention modules, transformers, and BERT. The concluding section suggests some possible directions for future research on this topic.
@@ 1. Introduction
Up to about 12 years ago, commonly used information processing tools generally extracted a feature vector — a list of features along with their associated numerical weights — from the form in which the information was initially provided: structured or unstructured text, images, audio, video, etc. These features typically corresponded to human-understandable properties or attributes of the original information. We confine our discussion in this article to textual information only. For text, features would include words and phrases occurring in the text, their counts, morphology, parts of speech, and various grammatical relationships. These features were chosen on the basis of linguistic considerations; they were, in a sense, interpretable by design. A number of studies (some of which are discussed in more detail in Section 6) did focus on understanding feature weights, their behaviour, and the mathematical / statistical models underlying their computation and use, but they were not explicitly designated as explainability / intepretability studies.
Over the past few years, advances in Deep Machine Learning (ML) techniques — in particular, the potent combination of Large Language Models (LLMs) and Retrieval Augmented Generation (RAG) systems — have revolutionised the state of the art in textual information processing (as well as a number of other areas). The traditional paradigm of document ranking — represented by “ten blue links” — is increasingly being replaced by interactive and conversational forms of information access (fntir-conv-ir). These radical improvements have come at a cost: text is no longer simply represented in terms of easily interpretable features. Instead, early in the information processing pipeline, text is converted to so-called “dense” vectors in $\mathbb{R}^{n}$; the dimensions of this space are no longer easily understood. This representation is then processed by ML systems that have large and complex architectures, involving many parameters, often numbering in the billions. As a fallout, there has been a marked increase in the attention devoted to providing high-level, intuitive explanations of these models and methods.
Thus, this seems to be an appropriate time to present an organised summary of the major research efforts in the area of explainability in textual information processing and retrieval. Indeed, several such general surveys have already been published (sogaard2021explainable; posthoc-interpret-neural-nlp; 10.1145/3529755; explain-ir-avishek-anand-survey), in addition to more specialised surveys, such as (belinkov:2019:tacl; position-info-transformer). Our objective is to complement these existing surveys by describing more recent work in areas already covered there, and by focussing on areas not addressed by them. Before discussing the relation between this survey and the existing ones in Section 2, we first briefly describe how we collected the set of papers reviewed in this survey.
Scope of this survey
@@
The principal terms used in our search were variants of the keywords explanation (including explain*) and interpretability (viz., interpret*). These terms have been variously defined within the community. For example, according to lime, explaining a model may be defined as presenting textual or visual cues that provide qualitative understanding of the relationship between the components of an input instance (e.g. words in text), and the model’s output for that instance. doshivelez2017towards define interpretability more generally as simply “the ability to explain or to present in understandable terms to a human”. We adopt the broad sense conveyed by these definitions, and use the terms explainability and interpretability interchangeably in this survey.
@@
Our preliminary search revealed a substantial body of work on explainability in Recommender Systems and Computer Vision. However, we excluded papers from these areas unless they directly contributed to follow-up work in IR or NLP. While IR systems and Recommender Systems share similarities in their goal of ranking items to meet users’ information needs, often with or without explicit queries, this survey focuses specifically on the retrieval of text documents. As text-focused IR methods are deeply intertwined with NLP techniques, we concentrate on the explainability of IR and NLP to maintain this focus. We include papers published from 2004 to 2025 in this set. Some influential papers published long before 2004 were also added. From the longlisted set of papers, a subset of 60 papers was selected for more detailed discussion in this article. We also provide short overviews, insights of several additional works; however, these are not included in the reported count.
@@
Organisation of the survey:  The remainder of this survey is organised as follows. Section 2 lists existing surveys that cover explainability in textual IR and NLP, and describes the relation between them and the current article. Section 3 introduces the basic terminology used in the literature on explainability, and presents a schema for classifying explainability research. The studies summarised in the rest of the article are broadly organised using this schema. Section 4 reviews some well-known approaches to explainability that are most commonly used as baselines in recent work. Section 5 briefly describes the metrics that are currently commonly used to measure the quality of explanations. Even though metrics are crucial to measuring progress, it is worth noting that evaluation protocols for explainability are yet to achieve a satisfactorily mature status. Section 6 is devoted to a detailed discussion of explainable document / passage ranking, and is itself divided into subsections by topic. Within each of these sections and subsections, different studies have been discussed in chronological order. The expanding field of explainability for Retrieval Augmented Generation (RAG) Systems is discussed in a separate section (Section 7). We conclude in Section 8 with a few research questions that we believe merit further investigation.
For completeness, we also provide a set of appendices (B through E) that are intended as supplements to existing surveys (specifically those by posthoc-interpret-neural-nlp and 10.1145/3529755). These appendices cover recent work (not discussed in (posthoc-interpret-neural-nlp; 10.1145/3529755)) on explainability for individual elements of deep text processing pipelines, viz., word embeddings, sequence models, attention mechanisms, and transformer architectures.
@@ 2. Related Surveys
In a monograph,  jimmy-lin-book provide a comprehensive survey of neural IR models, covering topics from multi-stage ranking architectures to dense retrieval techniques. The authors observe that IR models may be stratified into three fairly well-separated levels on the basis of effectiveness: transformer-based ranking strategies represent a definite advance over pre-transformer Neural Ranking Models (NRMs), which in turn improve significantly over “traditional” [^1] In this survey, we use this adjective to refer to all non-neural IR models, including relatively recently proposed ones. statistical ranking models. An intuitive (but precise) grasp of what leads to these quantum jumps in effectiveness seems eminently desirable. This, along with a number of other factors, has led to much interest in the explainability of NRMs. There has always been a substantial overlap between the IR and NLP communities. With the recent success of neural methods, the boundary between the disciplines has been blurred even further. Not surprisingly, therefore, research in explainable IR (ExIR) is significantly influenced by similar research in NLP, which in turn draws heavily from explainability studies in ML in general.
@@
Our search for a comprehensive summary of research on this broad subject led us to two articles that provide an overview of attempts to explain neural models for NLP (posthoc-interpret-neural-nlp; 10.1145/3529755). While posthoc-interpret-neural-nlp focus on post-hoc [^2] In Section 3, we provide a review of this and other basic terms commonly used in the context of interpretability / explainability. methods for interpreting neural NLP methods, the main emphasis in (10.1145/3529755) is on the explainability of word embedding vectors, recurrent neural networks (RNNs) and attention mechanisms.
@@
This being an area of active research, a non-trivial body of work exists that has not been covered by the above surveys, either because the work was published later, or because its topic was outside the specific scope of the surveys. In particular, until very recently, there were no similar overviews addressing ExIR in detail. The current study started as an attempt to address that lacuna. explain-ir-avishek-anand-survey have published an article focused on this specific topic. There is, inevitably, a good deal of overlap between explain-ir-avishek-anand-survey’s survey and Section 6 of this article, but we also cover a fair amount of work that is not discussed by them. Further, since our survey was prepared independently, we believe that it provides a somewhat different perspective on ExIR. Finally, given that ExIR often draws inspiration from explainability for NLP, we review (in Sections B through E) recent work on Explainable NLP that was not covered by the earlier surveys (posthoc-interpret-neural-nlp; 10.1145/3529755). In the rest of this section, we briefly discuss other related surveys that are either older, or have a scope that does not substantially overlap with ours.
@@
To the best of our knowledge, doshivelez2017towards were the first to attempt to formalise various notions related to interpretability in ML. lipton-16 provides an exposition of the broader aspects of interpretability in ML and Artificial Intelligence (AI), and why it is challenging, especially in the modern scenario. sogaard2021explainable examines extant taxonomies of explainability methods, and carefully constructs a framework for organising them. The next section draws from these articles to build a glossary of common terms used in this area.
@@
belinkov:2019:tacl cover explainable NLP, and summarise studies that analyse neural network nodes, and activations across different layers of deep neural networks, in order to understand what linguistic features are encoded by neural NLP models. As mentioned above, many of the discussed methods were based on general interpretability approaches proposed by the ML community, often within the context of Computer Vision / Image Processing problems. Signalling the growing interest in this area, danilevsky-etal-2020-survey also provide an early overview of explainability for NLP, describing different explanation categories and approaches, as well as several visualization techniques. 10.1162/tacl_a_00440 focus on explanation as a tool for debugging NLP models, with humans in the loop.
@@
As Transformer and BERT-based techniques were widely adopted in the IR / NLP community, several researchers focused on various aspects of BERT, from its knowledge-representation power to fine-tuning, pre-training, and over-parametrisation issues. rogers-etal-2020-primer were the first to survey research in this category, and present a summary of “what we know about how BERT works.” The surveys discussed at the beginning of this section provide a more up-to-date summary of research in this broad area.
@@
There are other broader efforts in Explainable AI (XAI), such as XAIR (Clement2023XAIRAS), which systematically reviews XAI methods aligned with the software development process. However, this meta-review targets general AI explainability rather than the specific challenges of textual and retrieval systems addressed in our survey.
@@ 3. Categorising Explanation Methods
Some ML models, e.g., rule-based systems, decision trees and sparse linear models, are regarded as inherently explainable, because it is relatively simple for a human to trace and describe their inner workings. Traditional IR models like BM25 are also generally regarded as inherently explainable. Such models comprise the class $\mathcal{S}$ (see Table 1). Explanations of such models are, naturally, model-aware. The composition of $\mathcal{S}$ is open to debate, however. pmlr-v119-moshkovitz20a argue that even apparently well-understood methods with a single parameter, viz., $k$-means and $k$-medians clustering techniques, may be regarded as “opaque”, because these methods define cluster boundaries on the basis of linear combinations of features, rather than a single feature value. More pertinent to our context is the debate on whether the attention mechanism can be construed as providing explanations. Section D.1 discusses this debate in more detail.
In contrast, post-hoc explanations primarily seek to elucidate the output, rather than the inner workings of models. Visualisations of word embeddings (e.g., using t-SNE (van2008visualizing)) are a form of post-hoc “explanation”: they do not explain why certain words are mapped to certain vectors, but once the vectors are generated (i.e., after the fact), they provide insights about the correspondence between proximity in the semantic and vector spaces. Explanations by example (e.g., “Model $M$ predicts class $Y$ for input instance $X$ because $X$ is ‘similar’ to training instances ${X}_{1},{X}_{2},\ldots,{X}_{k}$ which were labelled YY”) also fall within this category. Post-hoc explanations may be either model-aware or model-agnostic.
@@
Global methods explain the overall behaviour of a model. Global explanations should apply to all (or most) input-output pairs for a given model. An example of a global explanation in the IR context is: “ranking based on query likelihood tends to be biased in favour of short documents when Jelinek-Mercer smoothing is used instead of Dirichlet smoothing” (smucker2005investigation). On the other hand, a local explanation only seeks to explain a model’s decision for a given input instance.
Pairwise vs. pointwise vs. listwise local explanations
@@ 3.4.2. Feature attribution and rationales
Instead of explaining $\mathcal{M}_{C}$’s behaviour in terms of $\mathcal{M}_{S}$, researchers have also attempted to understand $\mathcal{M}_{C}$ by identifying which features of an input instance $I$ (i.e., which components of $I_{S}$ or $I_{C}$) most strongly affect $\mathcal{M}_{C}$’s output for $I$. This approach is called feature attribution. Several techniques for quantifying and visualising the contribution of a particular feature towards the prediction made by a model have been proposed in the literature. These local explanations may be provided graphically in the form of a heatmap or saliency map (saliency-map-simonyanVZ13), with different colours or shades representing different degrees of importance of features, or simply as a list of the most important features along with their importance scores.
A related approach involves the construction of a rationale, a concise representation (or the extracted essence) of $I$ that is (i) both short and simple enough to be readily understandable by a human, but (ii) for which $\mathcal{M}_{C}$’s prediction is the same as that for $I$. For example, a short excerpt from an input text that is sufficient to correctly predict the sentiment of a sentence comprises a rationale for a sentiment classification model (deyoung-etal-2020-eraser).
3.4.3. Contrastive explanations
Contrastive methods (jacovi-etal-2021-contrastive) aim to elucidate a model’s prediction for a particular instance $I$, by contrasting it with an alternative, closely related instance $I^{\prime}$ for which the prediction is different. These methods are, by design, local in nature. In practice, researchers (e.g., counterfactual-himabindu) frequently operationalise contrastive explanations through the use of counterfactuals, by identifying minimal perturbations to the input that result in a change in the model’s prediction. These minimal changes serve as the explanatory evidence, isolating the features most responsible for the observed decision. Contrastive explanations may thus also be regarded as a special case of feature attribution.
A possible instantiation of this idea within a ranking scenario would lead to pairwise document explanations (see Section 3.3). For example, if document $D_{i}$ is ranked higher than document $D_{j}$, a contrastive explanation might identify a set of “good” terms that contribute positively to the rank of $D_{i}$ because of their significant presence in $D_{i}$; these terms would typically be absent (or only weakly present) in $D_{j}$, leading to its inferior rank.
counterfactual-editing present a concrete application of counterfactual reasoning in a retrieval setup. Their method aims to minimally perturb the query, instead of the document. Given a query $Q$ and documents $D_{i},D_{j}$ (where $D_{i}$ is ranked higher than $D_{j}$), their objective is to construct a $Q^{\prime}$ by counterfactually editing $Q$, for which $D_{j}$ would be ranked higher than $D_{i}$. Specifically, their method first identifies important query terms using ColBERT term-weighting (Equation 1). The selected token is then masked and appended to the beginning of the counterfactual document ($D_{j}$). Next, DistilBERT is employed to predict replacement tokens, maintaining a beam of possible candidates. For each candidate, if the resulting retrieval score of $D_{j}$ surpasses that of $D_{i}$, the modified token is identified as the effective counterfactual edit; otherwise, additional tokens are iteratively modified until the ranking change is achieved. The proposed framework shows success on QA datasets, such as, MSMARCO, NaturalQA, FEVER, etc. While this study represents a first attempt, it has some drawbacks. There is no constraint on the maximum number of query terms that can be modified. Thus, the edited query may actually be entirely different from the original. Although such extensive edits might successfully promote the counterfactual document in the ranking, they risk altering or diminishing the original information need of the user.
attacks-sigir-23 and related studies ((prada-tois-23; nrm-robust-tois-2023)) also use similar ideas (involving document edits instead of query edits), but their focus is on studying the robustness of NRMs under such ‘attacks’, rather than explanations, per se. These articles are therefore not discussed in detail in our survey.